OBSERVATION OF HIGH-AMPLITUDE ZERO-POINT MOTION AND LOW-BARRIER HYDROGEN BONDS IN THE H_{3}O_{2}^{−} AND H_{2}O_{2}^{−} COMPLEXES USING ARGON PREDISSOCIATION INFRARED SPECTROSCOPY

\underline{ERIC G. DIKEN}, JEFFREY M. HEADRICK, JOSEPH R. ROSCIOLI, AND MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

We used argon predissociation infrared spectroscopy in the 600-3800 cm^{−}^{1} region to investigate the quantum nature of the shared proton in the [HO-H-OH]^{−} and the [O-H-OH]^{−} complexes. High-level theoretical calculations anticipate the zero-point levels to lie above the barriers in their respective proton transfer coordinates, consequently giving rise to highamplitude proton motion. Experimentally, very intense, sharp bands are observed below 800 cm^{−}^{1} for each complex, and are assigned to the fundamental (1←0) transitions due to the shared proton excitation. Single, weak features were also recovered in the 3650-3660 cm^{−}^{1} region for each complex, and are assigned to ”hydroxide-like” OH stretches that result from large-amplitude proton oscillation along the heavy atom axis. Assignments of the three fundamental transitions associated with the three-dimensional confinement of the shared proton in the H_{3}O_{2}^{−} species are offered with the aid of diffusion Monte Carlo (DMC) calculations. In the H_{2}O_{2}^{−} case a preliminary analysis of the O-O stretch contribution to the zero-point energy level is proposed.